8. The effects of property rights on achieving efficiency Consider a river found in the city of Philadelphia, and then answerComplete the following table by computing the total profit (the fish cannerys economic profit and the campgrounds economicfill in the blanks)

1) "production method is 'not used/used"

2) "the fish cannery will 'not use/use"

3) "pay the campground '0$/between 0-300$/between 300-700$/between 700-1000$"

4) "no cost, the fish cannery will 'not use/use"

5) "the most economic profit when ' the fish cannery has property rights to pollute the river/ it has property rights to a clean river."
Answer #1

No Recycling = 2700+1900 = 4600

Recycling = 2000+2900 = 4900

Blanks-

1) used

2) use

3) $0

4) use

5) between $700 and $1000

6) it has property rights to a pollute or clean the river

The statement is False

> All Correct. Read the question carefully.

Arun K Sun, Nov 14, 2021 11:24 PM